OpenStudy (anonymous):

For questions 1 through 6 graph each exponential function. f(x) = 6x(1 point) f(x) = (1 point) f(x) = 4x + 2(1 point) f(x) = (1 point) f(x) = 3x+1(1 points) f(x) =-3(1 points) Using complete sentences, explain the limitations on the base of an exponential function. (2 points) Using complete sentences, explain how the graphs of exponential functions are different from the other functions you’ve studied (e.g., linear, quadratic, cubic, rational, etc.)(2 points)

OpenStudy (anonymous):

please help

OpenStudy (jdoe0001):

I don't see any exponents there

OpenStudy (anonymous):

@jdoe0001 6^x for 2.( 1/7)^x 3.graph the function f(x) =f(x) = 4^(x + 2) 4.(1/3)^(x-3) 5.f(x) = 3^x +1 6.(1/2)^x -3

OpenStudy (jdoe0001):

well, then just pick a few values for "x", starting with say 0, and go negative and positive from there you'd get some "y" value for the function, so plot that for example on the 1st one $$\bf y= 6^x\\ x =0 \implies y = 6^o = 1\\ \color{blue}{\text{so }(0, 1)}\\ x =-2 \implies y = 6^{-2} = \frac{1}{6^2} = \frac{1}{36}\\ \color{blue}{\text{so }(-2, \frac{1}{36})}\\ x =+2 \implies y = 6^2 = 36\\ \color{blue}{\text{so }(2, 36} $$

OpenStudy (anonymous):

so then how would i graph it? cause my graph i attepted looks like this  @jdoe0001